# fftvis: A Non-Uniform Fast Fourier Transform Based Interferometric Visibility SimulatorA

### v1

@@ 2.5 Caveats and Limitations
The fftvis algorithm, as detailed in this section, provides an efficient NUFFT-based framework for simulating radio interferometric visibilities. The current implementation prioritizes performance for common 21 cm cosmology use cases, particularly drift-scan observations with identical antenna beams. While powerful within this scope, we highlight several areas present opportunities for future development to broaden its applicability.
@@ 3 The fftvis Package
The primary interface, fftvis.simulate_vis, accepts a sky model (as either a list of point sources or a HEALPix map), an array configuration, a beam model (supporting both analytic and gridded forms via the pyuvdata.BeamInterface), and metadata specifying the observing cadence and frequency channels. Internally, the simulator rotates the sky to topocentric coordinates at each time step using numpy matrix multiplication functions, evaluates the beam response with scipy interpolation routines, and grids the product onto the Fourier plane using finufft’s Type 3 NUFFT algorithm. The coordinate rotation code used in fftvis is adapted directly from the matvis codebase, ensuring consistent and efficient transformations and reducing the possibility of mismatched assumptions between simulators. This design choice further allows for the interoperability between the two tools. Parallelization over frequencies and time integrations is achieved via the ray library and native OpenMP threading within finufft.
@@ 4 Simulator Validation
[…]
@@
The first test focuses on evaluating the intrinsic numerical accuracy of fftvis by comparing its outputs against analytic visibility solutions derived from simplified diffuse sky models. These models admit closed-form expressions for the visibilities, allowing for an objective and implementation-independent benchmark (Lanman et al., 2022). The agreement with these solutions provides a strong validation of the accuracy of the mathematical and numerical implementation of the simulation. The second test assesses fftvis in the context of its intended use as a validator for 21 cm analysis pipelines. Specifically, we examine the spectral and temporal coherence of the simulated visibilities by computing their delay and fringe-rate spectra and comparing with the visibilities produced by the matvis simulator. With foregrounds exceeding the cosmological signal by 4–6 orders of magnitude, even subtle spectral artifacts from simulations can introduce spurious structure that contaminates the EoR window. To be useful as a validation tool, the simulator must avoid such artifacts and preserve the expected spectral smoothness of the simulated visibilities. This test examines whether fftvis achieves the dynamic range and fidelity necessary to support robust pipeline validation and foreground modeling. For all validation tests presented in this section, we use the following software versions: matvis=1.2.1, pyuvsim=3.0.0, and fftvis=1.0.0. Unless otherwise stated, we set the NUFFT precision to $\epsilon=10^{-13}$ for all tests.
@@ 4.1 Comparisons to Analytic Diffuse Models
Lanman et al. (2022) developed a suite of analytically tractable sky brightness distributions designed to test and validate the numerical accuracy of visibility simulators under realistic wide-field conditions. These test patterns are constructed to model the structural and angular characteristics of typical astrophysical emission (i.e., horizon-spanning profiles, sharp horizon cuts, etc.), while remaining mathematically simple enough to permit either closed-form or rapidly converging series solutions to the RIME. These analytic solutions provide an absolute reference against which the fundamental accuracy of a simulator’s implementation can be assessed. For our analysis, we consider three test cases:
@@
These models are discretized onto a full-sky HEALPix grid (Górski et al., 2005) with $N_{\mathrm{side}}=1024$ using the analytic_diffuse package [^7] https://github.com/aelanman/analytic_diffuse, and fed into both fftvis and matvis for evaluation. Including matvis in this comparison serves two key purposes. First, it provides a direct cross-check against an established simulator that has been independently validated. Second, since both fftvis and matvis operate on the same discretized HEALPix input in this test, comparing their outputs helps isolate any differences arising specifically from the core visibility calculation method (i.e., fftvis’s NUFFT approach versus matvis’s matrix-based direct summation) from the errors inherent in the point-source approximation used for the input sky model.
@@
The remaining discrepancies between the simulator outputs (for both fftvis and matvis) and the analytic solutions arise primarily not from errors in the visibility calculation algorithms themselves, but from the approximations inherent in using a finite HEALPix grid to represent a continuous diffuse sky. As discussed in Lanman et al. 2022, the point source approximation, used by both matvis and fftvis, leads to integration errors, particularly near the horizon where geometric delays are largest, the fringe term oscillates most rapidly across a pixel, and sharp cut-offs in sky brightness or beam response occur. Even small misrepresentations in the boundaries of the pixels can result in significant amplitude deviations at high $u$. These effects are inherent to any simulator using the point source approximation on pixelized maps, and diminish with increasing angular resolution (i.e., higher $N_{\mathrm{side}}$). The excellent agreement between fftvis and matvis confirms that the NUFFT-based gridding used in fftvis introduces no measurable bias relative to matvis’s direct summation at the tested tolerance levels.
@@ 4.2 Delay and Fringe-Rate Spectrum Dynamic Range
Because astrophysical foregrounds are expected to exceed the 21 cm signal by 4 to 5 orders of magnitude, even low-level numerical artifacts introduced by the simulator can compromise this separation. A sensitive metric of such contamination is the delay spectrum, obtained by Fourier transforming visibilities along the frequency axis,
[…]
where $T(\nu)$ is a spectral taper (typically a Blackman-Harris taper for HERA) applied to suppress spectral sidelobes (Parsons et al., 2012). In this domain, spectrally smooth foregrounds are confined to low-delay modes, while the 21 cm signal appears at higher delays. The delay-space dynamic range, defined as the ratio of power at low delays to residual power at high delays, provides a stringent test of the spectral accuracy of the simulator.
@@
Our simulations for these delay-spectrum tests use a realistic foreground sky model composed of diffuse Galactic synchrotron emission derived from the Global Sky Model (GSM; Zheng et al. 2017) combined with a population of unresolved extragalactic point sources based on the GLEAM source-count distribution (Franzen et al., 2019). The sky model is then pixelized onto a full-sky HEALPix grid of size $N_{\mathrm{side}}=256$. This resolution was chosen to adequately sample the sky structure in the GSM and prevent significant fringe aliasing for the longest baselines considered in the simulation. We simulate observations for four representative baselines drawn from the HERA array configuration: two east-west (14.6 m, 146.0 m) and two north-south (25.3 m, 193.9 m), selected to sample a range of geometric delays and orientations. The instrument’s primary beam is modeled as an ideal frequency-dependent Airy pattern corresponding to HERA’s 14-meter dish diameter, and the simulated observatory location is fixed to that of HERA. Visibilities are computed across a 100–120 MHz band using 160 channels, corresponding to a channel resolution of approximately 122 kHz.
@@
While decreasing $\epsilon$ (i.e., tightening precision) improves accuracy, it correspondingly increases the computational cost of the NUFFT. This cost increase is primarily influenced by the NUFFT gridding kernel width, $w$, which scales approximately as $\log(1/\epsilon)$. Several internal finufft operations depend on $w$, including aspects of source gridding and the determination of the intermediate FFT grid size (as detailed in Section 2.2), making their runtime contributions $\epsilon$-dependent. However, because $w$ grows only logarithmically with $1/\epsilon$, even substantial changes in the requested precision (e.g., from $\epsilon=10^{-7}$ to $10^{-14}$, which roughly doubles $w$) lead to relatively modest increases in the runtime components that scale polynomially with $w$. For instance, a component scaling as $w^{d}$ would increase by a factor of approximately $2^{d}$ (e.g., a factor of 4 if $d=2$). For this reason, we set the default value of $\epsilon=10^{-13}$ in fftvis to prioritize delay-space dynamic range over a small increase in simulator efficiency. However, we set this as a tunable parameter in the fftvis API, allowing the user to increase their simulation runtime if delay-space dynamic range is not a strict requirement of their simulated visibilities.
Beyond the spectral characteristics tested by delay spectra, accurately simulating the temporal evolution of visibilities is essential for validating end-to-end analysis pipelines. Many 21 cm analysis techniques rely on the assumption that astrophysical signals evolve smoothly over time to distinguish them from temporally variable systematics (Rath et al., 2024; Pascua et al., 2024; Garsden et al., 2024; Charles et al., 2024). Therefore, simulated visibilities must replicate the expected time structure of the sky signal without introducing spurious variations from the simulation process itself. As with the spectral axis, a potential concern is that the NUFFT gridding step in fftvis, which interpolates in $uv$-space independently at each timestep, may introduce subtle time-dependent artifacts. These could arise as the source coordinates evolve with Earth rotation, especially during intervals of rapid beam-weighted sky change such as when bright sources set on the horizon.
@@
Figure 4: Fringe-rate spectra of simulated visibilities for three baselines with east-west projections of $14.6\,\rm m$, $292.0\,\rm m$, and $584.0\,\rm m$, generated using fftvis (using $\epsilon=10^{-13}$) and matvis. The top panel shows the normalized amplitude of the fringe-rate transformed visibilities, while the bottom panel presents the absolute residuals between the two simulators. Vertical dotted lines indicate the analytically expected maximum fringe rate, corresponding to the east-west projection associated with each baseline and latitude of the simulated array (Parsons et al., 2016). We find that the residuals between fftvis and matvis are below $10^{-13}$ for all baselines and fringe rates tested. The excellent agreement across all three baselines demonstrates that the use of the NUFFT in fftvis does not introduce significant time-dependent artifacts into the simulated visibilities.
@@ 5 Computational Considerations
Our timing tests were conducted at the NRAO New Mexico Array Science Center (NMASC) [^8] https://www.aoc.nrao.edu/computing/, on a full node with eight 2.6 GHz E5-2670 Xeon dual-core processors, with the number of active cores constrained at runtime for specific test cases. Individual runtime tests were measured using the line_profiler Python package.
@@ 5.1 Runtime
Figure 5: Here, we demonstrate how execution times (top row) of matvis and fftvis scale with three key simulation parameters: (left column) the number of sources in the sky model, (middle) the number of antennas in a densely packed array, and (right) maximum baseline length in wavelengths. We begin from a fiducial simulation that uses a HEALPix sky model of $N_{\rm side}=256\ (\approx 7.8\times 10^{5}\ \rm pixels)$, a densely-packaged hexagonal array of $N_{\rm ants}=261$, and 32 time integrations at a single frequency channel. In each plot, we vary one parameter while holding the other parameters fixed. In the bottom row, we show the ratio of execution times (matvis/fftvis) for both single-core and 4-core CPU runs. From these panels, we find that fftvis excels when simulating many-element, dense arrays, but struggles when the input array becomes more sparse in $uv$-coordinates, due to an under-utilization of the FFT of the input sky. However, an important exception occurs when antenna positions form a grid. In this scenario, a Type 1 NUFFT enables more efficient visibility simulation, as shown in the right-hand column.
Figure 6: Breakdown of the relative contributions of the dominant sources of runtime as a function of the number of sources used for simulation. As expected, the finufft dominates the total runtime, but is only a factor of a few times more than coordinate rotation and beam interpolation, and scales at the same rate (linearly with the number of sources).
@@
The left column of Figure 5 demonstrates how the runtime of both simulators scales with the number of sources, $N_{\rm sources}$. Both fftvis and matvis exhibit approximately linear scaling, $\mathcal{O}(N_{\rm sources})$. This shared linear dependence can be partially attributed to operations performed per source component, primarily coordinate rotation and beam interpolation. While these steps contribute to the total execution time, the primary difference in performance is the result of the difference in the core visibility calculation methods. matvis constructs and multiplies an antenna response matrix (computational cost scaling roughly as $\mathcal{O}(N_{\rm ant}^{\alpha}N_{\rm sources})$, where the factor $\alpha$ depends on the efficiency of the matrix multiplication given the antenna/source matrix size), while fftvis grids each source onto the Fourier plane for evaluation via a Type 3 NUFFT.
@@
It is also important to note here that for arrays possessing a regular grid structure (either natively Cartesian or transformable to one, as detailed in Section 2.2.1), fftvis has the capability to use a Type 1 NUFFT. Our timing results demonstrate that this approach is indeed much more efficient than the Type 3 transform for varying $u_{\rm max}$. This efficiency primarily arises from how the size of its uniform target grid ($N_{\rm unif}$ is determined. For a given gridded array, $N_{\rm unif}$ is set by the extent and resolution of the array’s gridded baseline coordinates ($\mathbf{b}_{\rm gridded}$). Consequently, $N_{\rm unif}$ remains constant irrespective of the observing frequency $\nu$, making it independent of $u_{\rm max}$. With both $N_{\rm sources}$ (fixed by the experiment) and $N_{\rm unif}$ (fixed by the array’s physical baseline grid) being constant, the runtime for fftvis using a Type 1 NUFFT becomes largely insensitive to changes in $u_{\rm max}$. This behavior contrasts sharply with the default Type 3 NUFFT, where the internal grid $N_{\rm grid}$ necessarily scales with $u_{\rm max}$ (via $S_{i}$ from Equation 5) and also with the sky extent $X_{i}$.
@@
SKA-LOW (200 MHz) $1.3\times 10^{8}$ $5.8\times 10^{5}$ $1.04\times 10^{3}$ $1.14\times 10^{5}$

### v2

@@ 2.5 Caveats and Limitations
The fftvis algorithm, as detailed in this section, provides an efficient NUFFT-based framework for simulating radio interferometric visibilities. The current implementation prioritizes performance for common 21 cm cosmology use cases, particularly drift-scan observations with identical antenna beams. While powerful within this scope, we highlight several areas that present opportunities for future development to broaden its applicability.
@@ 3 The fftvis Package
The primary interface, fftvis.simulate_vis, accepts a sky model (as either a list of point sources or a HEALPix map), an array configuration, a beam model (supporting both analytic and gridded forms via the pyuvdata.BeamInterface), and metadata specifying the observing cadence and frequency channels. Internally, the simulator rotates the sky to topocentric coordinates at each time step using numpy matrix multiplication functions, evaluates the beam response with scipy interpolation routines, and grids the product onto the Fourier plane using finufft’s Type 3 NUFFT algorithm. The coordinate rotation code used in fftvis is adapted directly from the matvis codebase, ensuring consistent and efficient transformations and reducing the possibility of mismatched assumptions between simulators. This design choice further allows for the interoperability between the two tools. Parallelization over frequencies and time integrations is achieved via the ray library (2017arXiv171205889M) and native OpenMP threading within finufft.
@@ 4 Simulator Validation
Figure 2: Validation of fftvis against analytic full-sky visibility solutions for diffuse emission patterns from 2022ApJS..259...22L. Each panel shows the magnitude of the visibilities $|V|$ for a different sky models varying in zenith angle (top curves), along with absolute value of the difference $|V_{1}-V_{2}|$ (bottom curves) between fftvis and the analytic result (blue), and between fftvis and matvis (pink). All simulations were performed using discretized sky models generated by the analytic_diffuse package. fftvis matches the analytic solutions to better than $\sim$$10^{-5}$ for most of the range of $u$ values simulated, and agrees with matvis to near machine precision, confirming the accuracy and reliability of the NUFFT-based method for simulating smooth, diffuse sky structure. We note that this sky model error is the result of HEALPix gridding errors, and will decrease with the size of the input sky model.
@@
The first test focuses on evaluating the intrinsic numerical accuracy of fftvis by comparing its outputs against analytic visibility solutions derived from simplified diffuse sky models. These models admit closed-form expressions for the visibilities, allowing for an objective and implementation-independent benchmark (2022ApJS..259...22L). The agreement with these solutions provides a strong validation of the accuracy of the mathematical and numerical implementation of the simulation. The second test assesses fftvis in the context of its intended use as a validator for 21 cm analysis pipelines. Specifically, we examine the spectral and temporal coherence of the simulated visibilities by computing their delay and fringe-rate spectra and comparing with the visibilities produced by the matvis simulator. With foregrounds exceeding the cosmological signal by 4–6 orders of magnitude, even subtle spectral artifacts from simulations can introduce spurious structure that contaminates the EoR window. To be useful as a validation tool, the simulator must avoid such artifacts and preserve the expected spectral smoothness of the simulated visibilities. This test examines whether fftvis achieves the dynamic range and fidelity necessary to support robust pipeline validation and foreground modeling. For all validation tests presented in this section, we use the following software versions: matvis=1.2.1, pyuvsim=3.0.0, and fftvis=1.0.0. Unless otherwise stated, we set the NUFFT precision to $\epsilon=10^{-13}$ for all tests.
@@ 4.1 Comparisons to Analytic Diffuse Models
2022ApJS..259...22L developed a suite of analytically tractable sky brightness distributions designed to test and validate the numerical accuracy of visibility simulators under realistic wide-field conditions. These test patterns are constructed to model the structural and angular characteristics of typical astrophysical emission (i.e., horizon-spanning profiles, sharp horizon cuts, etc.), while remaining mathematically simple enough to permit either closed-form or rapidly converging series solutions to the RIME. These analytic solutions provide an absolute reference against which the fundamental accuracy of a simulator’s implementation can be assessed. For our analysis, we consider three test cases:
@@
These models are discretized onto a full-sky HEALPix grid (2005ApJ...622..759G) with $N_{\mathrm{side}}=1024$ using the analytic_diffuse package [^4] https://github.com/aelanman/analytic_diffuse, and fed into both fftvis and matvis for evaluation. Including matvis in this comparison serves two key purposes. First, it provides a direct cross-check against an established simulator that has been independently validated. Second, since both fftvis and matvis operate on the same discretized HEALPix input in this test, comparing their outputs helps isolate any differences arising specifically from the core visibility calculation method (i.e., fftvis’s NUFFT approach versus matvis’s matrix-based direct summation) from the errors inherent in the point-source approximation used for the input sky model.
@@
The remaining discrepancies between the simulator outputs (for both fftvis and matvis) and the analytic solutions arise primarily not from errors in the visibility calculation algorithms themselves, but from the approximations inherent in using a finite HEALPix grid to represent a continuous diffuse sky. As discussed in 2022ApJS..259...22L, the point source approximation, used by both matvis and fftvis, leads to integration errors, particularly near the horizon where geometric delays are largest, the fringe term oscillates most rapidly across a pixel, and sharp cut-offs in sky brightness or beam response occur. Even small misrepresentations in the boundaries of the pixels can result in significant amplitude deviations at high $u$. These effects are inherent to any simulator using the point source approximation on pixelized maps, and diminish with increasing angular resolution (i.e., higher $N_{\mathrm{side}}$). The excellent agreement between fftvis and matvis confirms that the NUFFT-based gridding used in fftvis introduces no measurable bias relative to matvis’s direct summation at the tested tolerance levels.
@@ 4.2 Delay and Fringe-Rate Spectrum Dynamic Range
A sensitive metric of such spurious spectral structure is the delay spectrum, obtained by Fourier transforming visibilities along the frequency axis,
$V_{ij}(\tau)=\int T(\nu)V_{ij}(\nu)e^{2\pi i\nu\tau}d\nu,$ (11)
where $T(\nu)$ is a spectral taper (typically a Blackman-Harris taper for HERA) applied to suppress spectral sidelobes (2012ApJ...756..165P). In this domain, spectrally smooth foregrounds are confined to low-delay modes, while the 21 cm signal appears at higher delays. The delay-space dynamic range, defined as the ratio of power at low delays to residual power at high delays, provides a stringent test of the spectral accuracy of the simulator.
@@
Our simulations for these delay-spectrum tests use a realistic foreground sky model composed of diffuse Galactic synchrotron emission derived from the Global Sky Model (GSM; 2017MNRAS.464.3486Z) combined with a population of unresolved extragalactic point sources based on the GLEAM source-count distribution (2019PASA...36....4F). The sky model is then pixelized onto a full-sky HEALPix grid of size $N_{\mathrm{side}}=256$. This resolution was chosen to adequately sample the sky structure in the GSM and prevent significant fringe aliasing for the longest baselines considered in the simulation. We simulate observations for four representative baselines drawn from the HERA array configuration: two east-west (14.6 m, 146.0 m) and two north-south (25.3 m, 193.9 m), selected to sample a range of geometric delays and orientations. The instrument’s primary beam is modeled as an ideal frequency-dependent Airy pattern corresponding to HERA’s 14-meter dish diameter, and the simulated observatory location is fixed to that of HERA. Visibilities are computed across a 100–120 MHz band using 160 channels, corresponding to a channel resolution of approximately 122 kHz.
@@
While decreasing $\epsilon$ (i.e., tightening precision) improves accuracy, it correspondingly increases the computational cost of the NUFFT. This cost increase is primarily influenced by the NUFFT gridding kernel width, $w$, which scales approximately as $\log(1/\epsilon)$. Several internal finufft operations depend on $w$, including aspects of source gridding and the determination of the intermediate FFT grid size (as detailed in Section 2.2), making their runtime contributions $\epsilon$-dependent. However, because $w$ grows only logarithmically with $1/\epsilon$, even substantial changes in the requested precision (e.g., from $\epsilon=10^{-7}$ to $10^{-14}$, which roughly doubles $w$) lead to relatively modest increases in the runtime components that scale polynomially with $w$. For instance, a component scaling as $w^{d}$ would increase by a factor of approximately $2^{d}$ (e.g., a factor of 4 if $d=2$). For this reason, we set the default value of $\epsilon=10^{-13}$ in fftvis to prioritize delay-space dynamic range over a small increase in simulator efficiency. However, we set this as a tunable parameter in the fftvis API, allowing the user to decrease their simulation run time if delay-space dynamic range is not a strict requirement of their simulated visibilities.
Beyond the spectral characteristics tested by delay spectra, accurately simulating the temporal evolution of visibilities is essential for validating end-to-end analysis pipelines. Many 21 cm analysis techniques rely on the assumption that astrophysical signals evolve smoothly over time to distinguish them from temporally variable systematics (2024arXiv240608549R; 2024arXiv241001872P; 2024MNRAS.535.3218G; 2024MNRAS.534.3349C). Therefore, simulated visibilities must replicate the expected time structure of the sky signal without introducing spurious variations from the simulation process itself. As with the spectral axis, a potential concern is that the NUFFT gridding step in fftvis, which interpolates in $uv$-space independently at each timestep, may introduce subtle time-dependent artifacts. These could arise as the source coordinates evolve with Earth rotation, especially during intervals of rapid beam-weighted sky change such as when bright sources set on the horizon.
@@
Figure 4: Fringe-rate spectra of simulated visibilities for three baselines with east-west projections of $14.6\,\rm m$, $292.0\,\rm m$, and $584.0\,\rm m$, generated using fftvis (using $\epsilon=10^{-13}$) and matvis. The top panel shows the normalized amplitude of the fringe-rate transformed visibilities, while the bottom panel presents the absolute residuals between the two simulators. Vertical dotted lines indicate the analytically expected maximum fringe rate, corresponding to the east-west projection associated with each baseline and latitude of the simulated array (2016ApJ...820...51P) where the slight bleed outside of this window is caused by extra temporal structure introduced by sources moving through the beam. We find that the residuals between fftvis and matvis are below $10^{-13}$ for all baselines and fringe rates tested. The excellent agreement across all three baselines demonstrates that the use of the NUFFT in fftvis does not introduce significant time-dependent artifacts into the simulated visibilities.
@@ 5 Computational Considerations
Our timing tests were conducted at the NRAO New Mexico Array Science Center (NMASC) [^5] https://www.aoc.nrao.edu/computing/, on a full node with eight 2.6 GHz E5-2670 Xeon dual-core processors, with the number of active cores constrained at runtime for specific test cases. Individual runtime tests were measured using the line_profiler Python package (line_profiler).
5.1 Review of the matvis algorithm
In order to contextualize the performance comparison between matvis and fftvis, we briefly summarize the matvis algorithm and its computational characteristics. The matvis simulator reformulates the RIME in a matrix-based framework that exploits the separability of per-antenna and per-source quantities. Rather than directly evaluating the full per-baseline sum, an operation that scales as $\mathcal{O}(N_{\rm sources}N_{\rm bls})$, matvis takes advantage of the correlation form of the measurement equation, in which a visibility can be expressed as a cross-correlation of antenna responses. This redefines the problem as $\mathcal{O}(N_{\rm sources}N_{\rm ants}^{\alpha})$, where $\alpha\approx 1.3-2$ depending on data layout and caching efficiency, offering a substantial reduction in runtime for large arrays compared to direct evaluators of the measurement equation, such as pyuvsim.
Starting from the interferometric measurement equation, the geometric phase factor for source $n$ observed by the baseline between antennas $i$ and $j$ can be separated into per-antenna terms:
$e^{-2\pi i\,\mathbf{b}_{ij}\!\cdot\!\mathbf{y}_{n}(t)}$ $=e^{-2\pi i\,\mathbf{x}_{i}\!\cdot\!\mathbf{y}_{n}(t)}\,e^{+2\pi i\,\mathbf{x}_{j}\!\cdot\!\mathbf{y}_{n}(t)}$
$\equiv\mathbf{F}_{in}\,\mathbf{F}_{jn}^{*},$ (12)
where $\mathbf{x}_{i}$ is the antenna position vector, $\mathbf{y}_{n}(t)$ encodes the direction cosine of source $n$ at time $t$, and $\mathbf{F}_{in}$ is the per-antenna phasor response. The intrinsic source coherency matrix $\mathbf{C}_{n}$ (which couples Stokes parameters to the instrumental polarization basis) can be factored as
$\mathbf{C}_{n}=\mathbf{M}_{n}\,\mathbf{M}_{n}^{\dagger},$ (13)
where $\mathbf{M}_{n}$ is a Cholesky-like decomposition convenient for vectorized computation. Combining the direction-dependent primary beam response $\mathbf{A}_{ipkn}$ with the phasor and coherency matrices, the visibility for polarization products $p,q$ is
$V^{pq}_{ij}(\nu_{a},t)=\sum_{n}\mathbf{A}_{ipkn}\,\mathbf{F}_{in}\,\mathbf{M}_{kk^{\prime\prime}n}\,\mathbf{M}_{k^{\prime}k^{\prime\prime}n}^{*}\,\mathbf{F}_{jn}^{*}\,\mathbf{A}_{jqk^{\prime}n}^{*}.$ (14)
To exploit matrix algebra, the per-antenna response to the entire sky is collected into an intermediate matrix
$\mathbf{Z}_{ip,\,nk^{\prime\prime}}=\sum_{k}\mathbf{A}_{ipkn}\,\mathbf{F}_{in}\,\mathbf{M}_{kk^{\prime\prime}n},$ (15)
which has shape $(N_{\rm ant}\!\times\!N_{\rm src})$ per polarization. The full visibility matrix is then obtained through an outer product:
$V^{pq}_{ij}(\nu_{a},t)=\left[\mathbf{Z}\,\mathbf{Z}^{\dagger}\right]_{ij}^{pq}.$ (16)
In this formulation, each column of $\mathbf{Z}$ encodes the complex beam-weighted sky response of a single antenna across all sources, while the matrix product $\mathbf{Z}\mathbf{Z}^{\dagger}$ efficiently computes all cross-correlations simultaneously. This converts the explicit per-baseline summation into a pair of dense matrix multiplications, achieving substantial reuse of per-antenna terms at the cost of storing the full $\mathbf{Z}$ matrix in memory whose size scales as $\mathcal{O}(N_{\rm sources}N_{\rm ants})$.
@@ 5.2 Runtime
Figure 5: Here, we demonstrate how execution times (top row) of matvis and fftvis scale with three key simulation parameters: (left column) the number of sources in the sky model, (middle) the number of antennas in a densely packed array, and (right) maximum baseline length in wavelengths. We begin from a fiducial simulation that uses a HEALPix sky model of $N_{\rm side}=256\ (\approx 7.8\times 10^{5}\ \rm pixels)$, a densely-packaged hexagonal array of $N_{\rm ants}=261$, and 32 time integrations at a single frequency channel. In each plot, we vary one parameter while holding the other parameters fixed. In the bottom row, we show the ratio of execution times (matvis/fftvis) for both single-core and 4-core CPU runs. From these panels, we find that fftvis excels when simulating many-element, dense arrays, but struggles when the input array becomes more sparse in uvuv-coordinates, due to an under-utilization of the FFT of the input sky. However, an important exception occurs when antenna positions form a grid. In this scenario, a Type 1 NUFFT enables more efficient visibility simulation, as shown in the right-hand column.
Figure 6: Breakdown of the relative contributions of the dominant sources of runtime as a function of the number of sources used for simulation for both matvis and fftvis. Here the computational steps that are shared between fftvis and matvis (coordinate rotation and beam interpolation) are marked with dashed colored lines while steps specific to each software package are marked by solid lines. As expected, the finufft stage dominates the total runtime for fftvis, but is only a factor of a few times more than coordinate rotation and beam interpolation, and scales at the same rate (linearly with the number of sources). For matvis, the total time is dominated by the final matrix multiplication step in Equation 16 and is typically many times more expensive than the coordinate rotation and beam interpolation stages of the algorithm.
@@
The left column of Figure 5 demonstrates how the runtime of both simulators scales with the number of sources, $N_{\rm sources}$. Both fftvis and matvis exhibit approximately linear scaling, $\mathcal{O}(N_{\rm sources})$. This shared linear dependence can be partially attributed to operations performed per source component, primarily coordinate rotation and beam interpolation. While these steps contribute to the total execution time, the primary difference in performance is the result of the difference in the core visibility calculation methods. matvis constructs and multiplies an antenna response matrix (computational cost scaling roughly as $\mathcal{O}(N_{\rm ant}^{\alpha}N_{\rm sources})$, where as discussed above, the factor $\alpha=1.3-2$ and depends on the efficiency of the matrix multiplication given the antenna/source matrix size), while fftvis grids each source onto the Fourier plane for evaluation via a Type 3 NUFFT.
@@
It is also important to note here that for arrays possessing a regular grid structure (either natively Cartesian or transformable to one, as detailed in Section 2.2.1), fftvis has the capability to use a Type 1 NUFFT. Our timing results demonstrate that this approach is indeed much more efficient than the Type 3 transform for varying $u_{\rm max}$. This efficiency primarily arises from how the size of its uniform target grid ($N_{\rm unif}$ is determined. For a given gridded array, $N_{\rm unif}$ is set by the extent and resolution of the array’s gridded baseline coordinates ($\mathbf{b}_{\rm gridded}$). Consequently, $N_{\rm unif}$ remains constant irrespective of the observing frequency $\nu$, making it independent of $u_{\rm max}$, as shown by the red line in Figure 5. With both $N_{\rm sources}$ (fixed by the experiment) and $N_{\rm unif}$ (fixed by the array’s physical baseline grid) being constant, the runtime for fftvis using a Type 1 NUFFT becomes largely insensitive to changes in $u_{\rm max}$. This behavior contrasts sharply with the default Type 3 NUFFT, where the internal grid $N_{\rm grid}$ necessarily scales with $u_{\rm max}$ (via $X_{i}$ from Equation 5) and also with the sky extent $S_{i}$.
@@
SKA-LOW (200 MHz) $1.3\times 10^{8}$ $5.8\times 10^{5}$ $04\times 10^{3} 1.1$$14\times 10^{5}$